OpenStudy (anonymous):

Solve 3y+7/8=5 Please explain if you can. If not-don't worry about it.

OpenStudy (jadeishere):

so it's asking what 'y' is. What you do to one side, you have to do to the other to isolate 'y' we have minus 7/8 from the left side of the equality sign, so we must do that to the other side, too. So 3y - 7/8 + 7/8 = 5 - 7/8 then 3y = 4 1/8 divide 4 and one-eighth by 3 Then you get the value of the variable :)